OpenStudy (anonymous):

FAN AND MEDAL!! Identify the vertex for the graph of y = -2x2 - 12x + 1. (-3, 19) (-3, 55) (3, -53) (3, -47)

OpenStudy (anonymous):

Please help im so stuck

OpenStudy (solomonzelman):

\(\large\color{black}{ y = -2x^2 - 12x + 1 }\) \(\large\color{black}{ y = -2(x^2 + 6x) + 1 }\) \(\large\color{black}{ y = -2(x^2 + 6x+9-9) + 1 }\) \(\large\color{black}{ y = -2(x^2 + 6x+9) +18+ 1 }\) \(\large\color{black}{ y = -2(x+3)^2 +19 }\)

OpenStudy (solomonzelman):

I made a mistake in my first reply.

OpenStudy (anonymous):

I want to say its b but I have a feeling its wrong

OpenStudy (solomonzelman):

why b?

OpenStudy (solomonzelman):

\(\large\color{black}{ y = a(x-h)^2 +k }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ y = -2(x-(-3)~)^2 +19 }\)

OpenStudy (anonymous):

Sorry my brother said it was b and then her tried explaining it to me but I think hes wrong... @SolomonZelman can you walk me through it?